OpenStudy (anonymous):

find the values of x where the extreme values of the function y=x^3-7x-5x occur.

OpenStudy (anonymous):

i hope this is calculus

OpenStudy (anonymous):

and i hope there is a typo there too check the question, because it looks funny

OpenStudy (anonymous):

could it be \( y=x^3-7x^2-5x\) ?

OpenStudy (anonymous):

no there is no typo its like this in the book and the answers choice i get is not there

OpenStudy (anonymous):

probably a typo in the book then why write \(x^3-7x-5x\) instead of \(x^3-12x\)

OpenStudy (anonymous):

\[y=x^3-7x-5x\]

OpenStudy (anonymous):

then start with \[y=x^3-12x\] making \[y'=3x^2-12\] and set that equal to zero and solve

OpenStudy (anonymous):

i get 2 and -2 but that is not my choices

OpenStudy (anonymous):

(a) x=-0.33, x=5 (b) x=0.33, x=-5 (c) x=0.33, x=5 (d) x=2, x=5 (e) x=0.65 ,x=7.65

OpenStudy (anonymous):

yeah so do i that is why i think it is a typo

OpenStudy (anonymous):

lets try \[y=x^3-7x^2-5x\\ y'=3x^2-14x-5\\ y'=(3x+1)(x-5)\]

OpenStudy (anonymous):

then you get critical points at \(x=-\frac{1}{3},x=5\)